# Learning to Optimize for Mixed-Integer Non-linear Programming

### v3

@@ C.2 Rounding with Straight-through Estimator
While this method does not involve any learning mechanism for rounding correction $\varphi$, the model in this ablation setup directly rounds to the nearest integer without leveraging information from the whole relaxed solution and problem coefficients. This helps to measure how learning-based rounding can refine the rounding process and enhance overall performance.
Appendix D Detailed Problem Setup and Parameter Sampling
D.1 Convex Quadratic and Simple Non-convex Problems
For the convex and non-convex quadratic problems, $\bm{Q}$ is a diagonal matrix with entries drawn from a uniform distribution over $[0,0.01]$ to guarantee the convexity of the problem. The $\bm{p}$ was generated with entries drawn from a uniform distribution over $[0,0.1]$. The matrix $\bm{A}$ was drawn from a normal distribution with a standard deviation of $0.1$. The parameter $\bm{b}\in\mathbb{R}^{m}$, the right-hand side vector of the inequality constraints, is sampled from a uniform distribution over the range $[-1,1]$.
D.2 Ronsenbrock Problems
For the Ronsenbrock Problems, the vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance and generated from a standard normal distribution.
The parameter $b$ is uniformly distributed over $[1,8]$ for each instance, and the parameter $\mathbf{a}\in\mathbb{R}^{n}$ represents a vector where elements drawn independently from a uniform distribution over $[0.5,4.5]$. The parameters $b$ and $\mathbf{a}$ influence the shape of the feasible region and the landscape of the objective function, serving as input features to the neural network.
Appendix E Experiments on Integer Linear Programs
In addition to the experiments on quadratic and non-convex problems described in the main text, we conducted further experiments on integer linear programming (ILP) problems. These experiments followed the same setup as the convex quadratic problems, except the quadratic terms were removed. Additionally, we imposed upper and lower bounds on the continuous decision variables, restricting $x$ to lie within the range $[-30,30]$:
$\min_{x\in\mathbb{Z}^{n}}\ \bm{p}^{\intercal}\bm{x}\;\text{subject to }\bm{A}\bm{x}\leq\bm{b},-30\leq\bm{x}\leq 30$
The ILP instances were generated with random coefficients for the objective function and constraint matrices in a manner consistent with the setup for the convex quadratic problems. Each instance featured varying right-hand side values, which were treated as input features. More details on the parameter generation process can be found in Appendix D.1.
Table 6: Result for an ILP Problem. Each problem size is evaluated on a test set of 100 instances. “Obj Mean” and “Obj Median” represent the mean and median objective values for this minimization problem, with smaller values being better. “Infeasible” denotes the fraction of infeasible solutions, and “Time (Sec)” is the average solving/inference time per instance. The “—” symbol indicates that no solution is found for any instance within 60 sec.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500
RC Obj Mean $0.477$ $-3.535$ $-13.335$ $-52.914$ $-71.006$ $-143.272$ $-344.111$
Obj Median $0.444$ $-3.094$ $-13.346$ $-53.119$ $-71.227$ $-143.775$ $-344.111$
% Infeasible $0\%$ $25\%$ $4\%$ $1\%$ $1\%$ $6\%$ $3\%$
Time (Sec) $0.0027$ $0.0028$ $0.0024$ $0.0034$ $0.0032$ $0.0034$ $0.0061$
LT Obj Mean $0.369$ $-1.244$ $-13.699$ $-52.786$ $-71.519$ $-150.723$ $-356.577$
Obj Median $0.354$ $-1.252$ $-13.721$ $-52.774$ $-71.520$ $-150.738$ $-356.575$
% Infeasible $4\%$ $24\%$ $0\%$ $0\%$ $0\%$ $3\%$ $7\%$
Time (Sec) $0.0029$ $0.0031$ $0.0028$ $0.0031$ $0.0034$ $0.0038$ $0.0063$
EX Obj Mean $-0.071$ $-6.543$ $-15.148$ $-56.341$ $-81.422$
Obj Median $-0.049$ $-7.096$ $-15.180$ $-56.374$ $-81.427$
% Infeasible $0\%$ $5\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.1303$ $0.1689$ $0.3773$ $61.2492$ $64.1954$
RR Obj Mean $-0.145$ $-6.706$ $-15.194$ $-56.505$ $-82.608$ $-172.916$ $-422.390$
Obj Median $-0.138$ $-7.227$ $-15.226$ $-56.511$ $-82.692$ $-172.927$ $-422.397$
% Infeasible $100\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$
Time (Sec) $0.1248$ $0.1289$ $0.1307$ $0.1461$ $0.1714$ $0.4149$ $1.9499$
N1 Obj Mean $2.533$ $-2.738$ $-3.110$ $-35.219$ $7.424$
Obj Median $2.563$ $-2.526$ $-2.971$ $-35.514$ $8.883$
% Infeasible $0\%$ $5\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.1241$ $0.1335$ $0.1362$ $0.1517$ $0.6259$
RL Obj Mean $0.158$ $-3.698$ $-14.004$ $-52.839$ $-74.392$ $-154.078$ $-379.467$
Obj Median $0.135$ $-3.429$ $-14.019$ $-52.862$ $-74.392$ $-154.061$ $-379.467$
% Infeasible $56\%$ $81\%$ $21\%$ $6\%$ $42\%$ $81\%$ $99\%$
Time (Sec) $0.0007$ $0.0008$ $0.0007$ $0.0009$ $0.0007$ $0.0009$ $0.0009$
Table 6 summarizes the results of the ILP experiments, covering problem sizes ranging from $5$ decision variables and $5$ constraints ($5\times 5$) up to $500\times 500$ with the same 60-second time limit. Overall, the learning-based methods (RC and LT) still demonstrated good performance in terms of feasibility and objective values compared to the baseline methods. In particular, the exact solver (EX, Gurobi) struggled with larger problem instances, while our methods maintained efficiency and scalability. This may be due to the random constraint matrices being dense.
However, an anomaly was observed for the learning-based methods in the $10\times 10$ case. Both RC and LT exhibited a relatively high proportion of infeasible solutions, with RC having 25% and LT having 24% infeasibility rates. This deviation stands out given the otherwise strong performance across other problem sizes, suggesting that these models may encounter difficulties when applied to some instances with specific parameter configurations. Generally, MILP solvers should be used for MILP problems. These experiments simply serve to show that our methods could be applied to MILPs.
Appendix F Training Time Comparison
@@
Table 10: Training Times (in seconds) for RC, LT, and RL methods across different problem sizes for the MILP Problem. Each method was set to train for 200 epochs, with early stopping applied.
Method 5x5 10x10 20x20 50x50 100x100 200x200 500x500
RC 213.35 274.80 115.68 169.82 104.95 73.87 149.78
LT 270.57 97.47 148.06 152.03 132.28 163.37 234.96
RL 114.89 68.97 70.58 69.90 86.46 54.87 144.26
Table 7, 8, 9 and 10, summarize the training times (in seconds) required by each method for problems of different scales. These results offer a clear view of the computational demands as problem complexity grows.

### v4

@@ Appendix F Ablation Study
Results and Insights.
The results of the ablation experiments, summarized in Table 2, demonstrate the importance of the correction layers $\varphi_{\Theta_{2}}$ in improving both solution quality and feasibility. RL shows a significant drop in feasibility rates, highlighting the importance of incorporating learnable corrective adjustments during training. Similarly, while RS benefits from differentiability via STE, the lack of learnable correction limits its performance compared to RC and LT.
Table 5: Ablation Study for Convex Quadratic Problems. See the caption of Table 2 for details.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500 1000×1000
RL Obj Mean $0.567$ $-2.325$ $-4.576$ $-13.932$ $-16.734$ $-35.562$ $-88.217$ $-170.738$
Obj Median $0.405$ $-2.654$ $-4.606$ $-13.952$ $-16.803$ $-35.545$ $-88.217$ $-170.738$
% Infeasible $24\%$ $74\%$ $23\%$ $33\%$ $62\%$ $73\%$ $100\%$ $100\%$
Time (Sec) $0.0004$ $0.0005$ $0.0004$ $0.0004$ $0.0006$ $0.0005$ $0.0005$ $0.0011$
RS Obj Mean $1.024$ $-0.973$ $-3.677$ $-11.542$ $-10.092$ $-22.366$ $-53.036$ $-105.194$
Obj Median $0.880$ $-1.269$ $-3.706$ $-11.525$ $-10.073$ $-22.483$ $-53.036$ $-105.194$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.0010$ $0.0010$ $0.0010$ $0.0011$ $0.0012$ $0.0012$ $0.0016$ $0.0032$
Table 6: Ablation Study for Simple Non-Convex Problems. See the caption of Table 3 for details.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500 1000×1000
RL Obj Mean $0.241$ $0.595$ $-0.138$ $-0.629$ $-1.581$ $-4.196$ $-11.531$ $-23.639$
Obj Median $0.090$ $0.484$ $-0.148$ $-0.655$ $-1.554$ $-4.196$ $-11.531$ $-23.639$
% Infeasible $28\%$ $46\%$ $13\%$ $49\%$ $85\%$ $100\%$ $100\%$ $100\%$
Time (Sec) $0.0005$ $0.0005$ $0.0005$ $0.0005$ $0.0006$ $0.0006$ $0.0006$ $0.0013$
RS Obj Mean $0.453$ $1.398$ $0.292$ $1.734$ $2.849$ $4.921$ $9.511$ $25.358$
Obj Median $0.328$ $0.983$ $0.284$ $1.736$ $2.841$ $4.907$ $9.511$ $25.358$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.0011$ $0.0011$ $0.0012$ $0.0011$ $0.0012$ $0.0013$ $0.0018$ $0.0031$
Table 7: Ablation Study for Rosenbrock Problems. See the caption of Table 4 for details.
Method Metric 2×4 20×4 200×4 2000×4 20000×4
RL Obj Mean $58.34$ $63.70$ $605.88$ $6221.75$ $68364.18$
Obj Median $58.00$ $61.95$ $609.00$ $5949.55$ $69087.00$
% Infeasible $86\%$ $36\%$ $44\%$ $28\%$ $31\%$
Time (Sec) $0.0006$ $0.0005$ $0.0005$ $0.0008$ $0.0014$
RS Obj Mean $25.09$ $69.36$ $684.67$ $6852.08$ $72910.55$
Obj Median $25.35$ $68.58$ $663.10$ $6509.01$ $68904.42$
% Infeasible $0\%$ $3\%$ $0\%$ $1\%$ $39\%$
Time (Sec) $0.0010$ $0.0010$ $0.0012$ $0.0019$ $0.0103$
Appendix G Details for Constraints Violations
In this section, we analyze constraint violations for three benchmark problems. The analysis focuses on both the frequency and magnitude of constraint violations, visualized through heatmaps for a comprehensive understanding. Each heatmap (Figure 6, Figure 7 and Figure 8) illustrates rows as test instances and columns as individual constraints.
Figure 6: Illustration of Constraint Violation Heatmap for 500×500 Convex Quadratic Problem for RC method (Top) and LT method (bottom) on 100 test instances: Each row represents an instance in the test set, while each column corresponds to a specific constraint.
Figure 7: Illustration of Constraint Violation Heatmap for 500×500 Simple Non-Convex Problem for RC method (Top) and LT method (bottom) on 100 test instances: Each row represents an instance in the test set, while each column corresponds to a specific constraint.
The heatmap for the convex quadratic problem (Figure 6) and the simple non-convex problem (Figure 7) reveals a sparse distribution of violations, predominantly concentrated in a single constraint. This indicates that most constraints are consistently satisfied, with only a few isolated violations. Notably, 4 out of 100 test instances of convex quadratic problem under RC are infeasible, and among these, 3 violations occur within the same constraint. Overall, the constraint violations are nearly negligible, confirming the effectiveness of the proposed methods.
Figure 8: Illustration of Constraint Violation Heatmap for 20000×4 Rosenbrock Problem for RC method (Top) and LT method (bottom) on 100 test instances: Each row represents an instance in the test set, while each column corresponds to a specific constraint.
Figure 8 highlights a much denser distribution of violations, reflecting the complexity of this benchmark. The nonlinear constraint $\|\mathbf{x}\|_{2}^{2}\leq nb$ is particularly challenging, as shown by the more frequent and larger violations.
The heatmaps reveal key insights into the performance of the RC and LT methods. While the convex quadratic and simple non-convex problems exhibit minimal violations, the Rosenbrock problem highlights the difficulty of satisfying nonlinear constraints. These observations underscore the need for further refinement of penalty weights. Specifically, constraint-specific adjustments could mitigate violations by placing higher penalties on constraints that are harder to satisfy.
Appendix H Experiments on Binary Linear Programs
Dataset.
For our experiments involving mixed-integer linear programs (MILPs), we utilized the ‘Obj Series 1‘ dataset from the MIP Workshop 2023 Computational Competition (Bolusani et al., 2023). This dataset comprises 50 related MILP instances derived from a common mathematical formulation, where the instances differ in a subset of the objective function coefficients. Each instance contains 360 binary variables and 55 constraints, with 120 out of the 360 objective coefficients varying across instances. All other components of the problem remain consistent.
Model Configuration.
The neural network architecture and hyperparameters were consistent with those used for other experiments in the main paper. Specifically for the MILP problem in this study, the input dimension of the neural network was set to 120, corresponding to the number of varying objective function coefficients, and the output dimension was set to 360, representing the binary decision variables. The hidden layer consisted of 256 neurons.
Results.
Table 8 summarizes the results of the ILP experiments: Both learning-based methods (RC and LT) demonstrate the ability to generate high-quality feasible solutions efficiently, with RC even surpassing the heuristic-based method N1 in terms of objective value. However, N1 is the fastest method overall, showcasing the robustness and efficiency of the heuristic in the MILP solver. Notably, the training time for the learning-based models is approximately 120 seconds, making them well-suited for applications requiring repeated problem-solving.
Table 8: Comparison of Optimization Methods on the MILP. See the caption of Table 2 for details.
Method Obj Mean Obj Median % Infeasible Time (Sec)
RC $9745.90$ $9763.00$ $0\%$ $0.04$
LT $14149.00$ $14149.00$ $0\%$ $0.04$
EX $8756.80$ $8747.00$ $0\%$ $289.13$
N1 $11901.10$ $11933.00$ $0\%$ $0.01$
Appendix I Training Time Comparison
@@
Table 9, 10 and 11, summarize the training times (in seconds) required by each method for problems of different scales. These results offer a clear view of the computational demands as problem complexity grows.